when molten magnesium oxide was electrolyzed for a certain period, 150 mg of Mg was deposited on the cathode. the volume of oxygen gas in cybic centimeter at STP conditions liberated at the anode during the same period is
1
Expert's answer
2020-12-18T08:41:49-0500

Solution:


At Cathode "-Mg^{+2}+2e^ - \;\implies Mg"

At Anode "-2O^{2-}\implies\;O_2\;+4e^-"


Overall reaction:

"Mg^{+2}+2O^{2-}\implies\;2Mg+O_2"


So, according to the reaction:


Since 2 moles of Mg releases 1 mole of O2 gas.


"\\dfrac{0.15}{24}\t\\;moles= \\dfrac{0.15}{48}\\;moles\\;of\\;O_2\\;gas= \\dfrac{0.15}{48}\t\t*22.4L=0.07L"


Answer:


"0.07L\;=70mm^3\; of \;O_2\;gas\;at\;STP\;."